OpenStudy (anonymous):

What single step would eliminate the x variable in the system of equations?    x − 2y = 7 3x − y = 15

OpenStudy (mertsj):

There is no single step that would do that.

OpenStudy (anonymous):

Multiply the top equation by 3 and subtract 3x-y = 15 from 3(x-2y = 7) so it would be 3x - 6y = 21 - (3x - y = 15) which should get you -5y = 6

OpenStudy (mertsj):

@asdfvg That is not a "single step".

OpenStudy (anonymous):

Then there isn't any other way haha

OpenStudy (anonymous):

a. Multiply the first equation by −3 b. Multiply the second equation by 7 c. Multiply the first equation by 2 d. Multiply the second equation by 5 Here are the options

OpenStudy (anonymous):

A

OpenStudy (doc.brown):

multiply the first row by -3 and add it to the second row

OpenStudy (anonymous):

didn't know there were options but none of those choices are technically in one step...you are multiplying by something and then you are adding it to the second equation to eliminate the x

OpenStudy (mertsj):

A

OpenStudy (doc.brown):

3x-3x-2y+5y=15-21

OpenStudy (doc.brown):

or more precisely, add -3 times the first row to the second